Fred has a shoebox of baseball cards. When he counted them by 2's, 3's, 4's, 5's, or 6's he had 1 left over. If he had fewer than 100 cards, how many did he have?
Answer by checkley77(12844)

The magic number is 61.
61/2=30 r1.
61/3=20 r1.
61/4=15 r1.
61/5=12 r1.
61/6=10 r1.